# An Elliott intertwining approach to classifying actions of $\mathrm{C}^{*}$-tensor categories

### 29 Jan 2024 v2

@@ Abstract.
We introduce a categorical approach to classifying actions of $*{}^{*}$-tensor categories $\mathcal{C}$ on $\mathrm{C}^{*}$-algebras up to cocycle conjugacy. We show that, in this category, inductive limits exist and there is a natural notion of approximate unitary equivalence. Then, we generalise classical Elliott intertwining results to the $\mathcal{C}$-equivariant case, in the same fashion as done by Szabó for the group equivariant case in [40].
@@ Introduction
Later, Popa classified amenable subfactors $N\subset\mathcal{R}$ by their standard invariant ([36]). Following subsequent reformulations ([29, 30]), the standard invariant of a finite index subfactor $N\subset\mathcal{R}$ can be understood as a pair $(F,Q)$ with $F$ an action of a unitary tensor category $\mathcal{C}$ on $N$ and $Q$ a special object in $\mathcal{C}$ that allows one to recover the inclusion $N\subset\mathcal{R}$ as a generalised crossed product. [^1] Precisely, $Q$ is a $Q$-system as defined by Longo in [29]. The general setting of $\mathrm{C}^{*}$-tensor categories represents a unifying framework for studying group actions and more general quantum symmetries arising from subfactor theory. Following the recent spectacular classification results for group actions on Kirchberg algebras by group equivariant KK-theory ([17]), and the development of a $\mathcal{C}$-equivariant KK-theory ([1]), it is natural to consider to what extent one can classify actions of more general $*{}^{*}$-tensor categories on simple, amenable $*{}^{*}$-algebras.
@@
Prior and closely linked to the classification of symmetries of operator algebras was the classification of the operator algebras themselves. First, Connes classified amenable factors in [8] with the exception of one type which was later completed by Haagerup ([22]). In the case of $*{}^{*}$-algebras, Elliott classified $*{}^{*}$-inductive limits of finite dimensional $*{}^{*}$-algebras by their ordered K-theory in [10]. Elliott’s methods laid the foundations for a roadmap to a classification of more general simple, amenable $*{}^{*}$-algebras. Following Elliott’s strategy and building on decades of work by many mathematicians, the classification programme of simple, amenable $\mathrm{C}^{*}$-algebras successfully culminated in [27, 35, 20, 21, 41, 5].
@@
The general strategy for a classification of operator algebras is to first achieve existence and uniqueness results for morphisms with respect to the proposed classification invariant. Loosely speaking, to construct an isomorphism between two objects $A$ and $B$ in the category of $\mathrm{C}^{*}$-algebras, by virtue of an existence type statement, one obtains morphisms $\phi:A\to B$ and $\psi:B\to A$ which induce mutually inverse elements at the level of the chosen abstract invariant. Then, a suitable uniqueness result will give that $\psi\circ\phi$ is equivalent to $\mathrm{id}_{A}$ and $\phi\circ\psi$ is equivalent to $\mathrm{id}_{B}$. In the category of separable $\mathrm{C}^{*}$-algebras, the required notion of equivalence is approximate unitary equivalence. The final step is to perform a so-called approximate intertwining argument. Essentially, one can tweak the morphisms $\phi$ and $\psi$ by unitaries until they become mutually inverse isomorphisms (see [38, Corollary 2.3.4] and the introduction of [40] for a more detailed breakdown of this roadmap to classification). Furthermore, in [12], Elliott proposes a similar classification strategy for arbitrary objects in more general categories that have a suitable notion of inner automorphisms.
@@
Instances of the implementation of Elliott’s strategy for the classification of compact (quantum) group actions on $*{}^{*}$-algebras appear in [18, 2], where the actions are assumed to have the Rokhlin property. In [40], Szabó articulates Elliott’s proposed strategy in the generality of actions of locally compact groups as an alternative to the Evans–Kishimoto intertwining type arguments (see [13]). In his work, Szabó defines an appropriate category of $\Gamma$$*{}^{*}$-algebras, that is $*{}^{*}$-algebras carrying an action of $\Gamma$, and produces a vast collection of intertwining results for this category. Szabó’s construction is a key ingredient facilitating the recent groundbreaking classification of amenable actions on Kirchberg algebras in [17].
@@
In this paper, we will develop the necessary techniques to perform approximate intertwining arguments for $*{}^{*}$-algebras carrying an action of a tensor category $\mathcal{C}$ ($\mathcal{C}$-$\mathrm{C}^{*}$-algebras). Glimpses of these techniques can be found in [6], where exact intertwining arguments are performed to classify inductive limit actions of fusion categories on AF-algebras. However, for classification, it is often necessary to develop a more general framework which allows to perform approximate intertwining arguments. In the setting of $\mathrm{C}^{*}$-algebras, Elliott used exact intertwining arguments to classify AF-algebras ([10]), but more refined approximate versions were needed to classify simple $A\mathbb{T}$-algebras in [11]. Approximate intertwining arguments in the setting of actions of tensor category have already appeared in the literature; see [42] for an adaptation of the Evans-Kishimoto intertwining argument introduced in [13].
@@
Primarily, the categorical framework developed by Szabó in [40] provides the conceptual skeleton for our construction. However, unlike in the group action case, a tensor category might not act by automorphisms, so our techniques differ. In [40], Szabó defines a cocycle morphism as a pair consisting of an extendible $*{}^{*}$-homomorphism and a unitary cocycle. In the setting of tensor categories, the action on a $\mathrm{C}^{*}$-algebra is given by a family of bimodules acting compatibly with respect to the tensor product. Therefore, the cocycles will be given by certain bimodule maps satisfying a family of commuting diagrams. Hence, it is not apparent how to adapt Szabó’s arguments.
[…]
@@
In $\mathrm{C}^{*}_{\mathcal{C}}$, we are able to construct inductive limits and define a suitable topology on the space of morphisms. An important challenge in constructing a category of cocycle morphisms is the composition. Composing two cocycle morphisms might not give a cocycle morphism. Therefore, we need to introduce a slightly different composition which will allow us to obtain a category. If the morphisms are non-degenerate, this new composition agrees with the canonical one, and with the one introduced by Szabó in the group action case ([40, Proposition 1.15]). Moreover, to avoid the need of restricting to non-degenerate morphisms, the maps $\mathbbm{v}_{X}$ are assumed to be isometries. This assumption allows us, unlike in [40], to work in the possibly non-extendible setting, generalising the notion of a cocycle morphism from [40] when restricting to twisted group actions.
@@
An essential observation for our construction is that we can encode the information of a cocycle morphism into a family of linear maps satisfying certain conditions (see [6, Lemma 3.8] where this is done for unital injective cocycle morphisms). Therefore, a cocycle morphism can be equivalently defined as follows.
@@
Let $\mathcal{C}$ be a $\mathrm{C}^{*}$-tensor category acting on $\mathrm{C}^{*}$-algebras $A$ and $B$ via $(A,F,J)$ and $(B,G,I)$ respectively. Then a cocycle morphism $(\phi,h):(A,F,J)\rightarrow(B,G,I)$ is given by a $*{}^{*}$-homomorphism $\phi:A\to B$ and a family of linear maps:
@@
    the following diagram commutes:
[…]
  5. (v)
[…]
In the case of a group action, we give an explicit formula for the family of linear maps $\{h^{X}\}_{X\in\mathcal{C}}$ in Example 3.14. Compared with the usual definition of a cocycle morphism given by pairs $(\phi,\mathbbm{v})$ consisting of a $*{}^{*}$-homomorphism and a family of bimodule maps $\{\mathbbm{v}_{X}\}_{X\in\mathcal{C}}$ (see Definition 3.1), Definition A provides a more direct framework for setting up the intertwining arguments. In particular, to compose two cocycle morphisms, one composes the $*{}^{*}$-homomorphisms, as well as the corresponding linear maps. Furthermore, the natural topology on cocycle morphisms is obtained by considering pointwise differences of the linear maps. When the acting category is semisimple and has countably many isomorphism classes of simple objects (which is assumed for the rest of the introduction), we can phrase convergence in this topology purely in terms of the linear maps.
@@
In general, this topology is coarser than the one used by Szabó in the group action case ([40, Definition 2.5]) (see Example 4.10). However, it is the same when restricted to non-degenerate cocycle morphisms. With this topology in hand, we can then define approximate unitary conjugation for cocycle morphisms. Moreover, unlike Szabó’s definition in [40, Definition 2.8], this is a symmetric relation (see Lemma 4.16), which simplifies the intertwining arguments. This can be formulated as follows.
@@
Along with Theorem D, in Section 5 we use our construction to obtain various $\mathcal{C}$-equivariant versions of classical intertwining arguments. For instance, we establish a general Elliott two-sided intertwining argument (see [38, Proposition 2.3.2] for a non-equivariant version) as well as an intertwining through reparametrisation in Theorem 5.9.
In Section 6, the techniques we develop allow us to perfom a $\mathcal{C}$-equivariant one-sided intertwining argument (see Theorem 6.1), generalising the classical one-sided intertwining arguments of [38, Proposition 2.3.5] and [40, Proposition 4.3]. This result is a key ingredient in the work of Evington, C. Jones, and the first named author in obtaining a McDuff-type characterisation of equivariant $\mathcal{D}$-stability of an action of a unitary tensor category, for strongly self-absorbing $\mathcal{D}$ ([15]). Precisely, $\mathcal{D}$-stability of an action is equivalent to the existence of a unital embedding of $\mathcal{D}$ into an appropriate subalgebra of the Kirchberg central sequence algebra of $A$ (in the case of a group action, this subalgebra coincides with the fixed point subalgebra of Kirchberg’s central sequence algebra). This generalises the classical result of [43, Theorem 2.2] (see also [27]) and its group equivariant counterpart (see [39, Corollary 3.8]). The remaining part of Section 6 is concerned with asymptotic versions of the results in Section 5.
@@ 1.1. Hilbert bimodules
If $X$ only satisfies the properties (i)-(iv) above, then we say $X$ is a pre-Hilbert $B$-module. For $x\in X$, we denote by $|x|^{2}$ the positive element of $B$ given by $\langle x,x\rangle_{B}$. To obtain a right-Hilbert $A$-$B$-bimodule, we need a compatible left action of a $\mathrm{C}^{*}$-algebra $A$.
Definition 1.2.
([25, Definition 2.1]) Let $A$ and $B$ be $\mathrm{C}^{*}$-algebras and $X={}_{A}X_{B}$ a bimodule over the complex algebras $A$ and $B$. We say that $X$ is a (right-)Hilbert $A$-$B$-bimodule if $X$ is a right-Hilbert $B$-module, and for all $a\in A$, the map $\phi(a):x\in X\mapsto ax\in X$ is adjointable, with adjoint $\phi(a)^{*}=\phi(a^{*})$.
In particular, $\phi:a\in A\to\phi(a)\in\mathcal{L}(X)$ is a $*{}^{*}$-homomorphism from $A$ to the $*{}^{*}$-algebra $\mathcal{L}(X)$ of adjointable maps on $X$. The map $\phi$ will be referred to as the left action of $A$ on $X$. For a Hilbert $A$-$B$-bimodule $X$, we often write the left action of an element $a\in A$ on a vector $x\in X$ by $a\rhd x$ and the right action of an element $b\in B$ on $x$ by $x\lhd b$.
@@
For a $*{}^{*}$-algebra $B$, the vector space $X=B$ is a Hilbert $B$-module with the right $B$-module structure given by multiplication and the inner product $\langle a,b\rangle_{B}=a^{*}b$ for $a,b\in B$. In this case, any $*{}^{*}$-homomorphism $\phi:A\rightarrow M(B)$ induces a right-Hilbert $A$-$B$-bimodule which we denote by ${}_{\phi}B$.
Let $A,B,C$ be $\mathrm{C}^{*}$-algebras. If $X$ is a Hilbert $A$-$B$-bimodule and $E$ is a Hilbert $B$-$C$-bimodule we may form their internal tensor product $X\boxtimes E$ that is an Hilbert $A$-$C$-bimodule. We sketch this construction and refer to [28, Section 4] for details. To perform the internal tensor product one starts by considering the algebraic tensor product of vector spaces $X\odot E$. We identify the elements of the form $x\lhd b\odot y$ with $x\odot b\rhd y$ to form the quotient
@@
$\langle x\boxtimes y,z\boxtimes w\rangle_{C}=\langle x,\langle y,z\rangle_{B}\rhd w\rangle_{C},$
for any $x,z\in X$ and $z,w\in E$. It follows that $V$ equipped with this $C$-action and $C$-valued inner product satisfies (i)-(iv) of Definition 1.1. We produce a Hilbert $C$-module $X\boxtimes E$ by completing $V$ under the norm defined by the inner product. Moreover, one can induce a left $A$ action on $X\boxtimes E$ through
@@
for all $a\in A$, $x\in X$ and $y\in E$. This equips $X\boxtimes E$ with the structure of a Hilbert $A$-$C$-bimodule.
If the bimodules are given by $*{}^{*}$-homomorphisms into the multiplier algebra as in Example 1.3, one may get greater insight into the structure of their tensor product. First we recall the definition of non-degenerate and extendible $*{}^{*}$-homomorphisms.
@@
Let $A$ and $B$ be $\mathrm{C}^{*}$-algebras and $\phi:A\to\mathcal{M}(B)$ a $*{}^{*}$-homomorphism. Then $\phi$ is said to be non-degenerate if $\phi(A)B$ is dense in $B$.
If $\phi:A\to\mathcal{M}(B)$ is non-degenerate, then $\phi$ extends to a unital $*{}^{*}$-homomorphism $\phi:\mathcal{M}(A)\to\mathcal{M}(B)$ ([28, Proposition 2.5]).
@@
Let $A$ and $B$ be $\mathrm{C}^{*}$-algebras. A $*{}^{*}$-homomorphism $\phi:A\to\mathcal{M}(B)$ is called extendible if for any increasing approximate unit $e_{\lambda}\in A$, the net $\phi(e_{\lambda})\in\mathcal{M}(B)$ converges strictly to a projection $p\in\mathcal{M}(B)$. [^5] Any non-degenerate $*{}^{*}$-homomorphism $\phi:A\rightarrow M(B)$ is extendible as for an approximate unit $e_{\lambda}\in A$ the net $\phi(e_{\lambda})$ converges strictly to $1_{\mathcal{M}(B)}$ (see e.g. [3, II.7.3.8.]). As discussed in [40, Definition 1.7], $\phi$ then factorises through $\mathcal{M}(pBp)\cong p\mathcal{M}(B)p\subseteq\mathcal{M}(B)$, with the $*{}^{*}$-homomorphism $\phi_{p}:A\to\mathcal{M}(pBp)$ now non-degenerate. In this case, following [40], we let $\phi^{\dagger}:\mathcal{U}(\mathcal{M}(A))\to\mathcal{U}(\mathcal{M}(B))$ be the unital group homomorphism defined by $\phi^{\dagger}(u)=\phi_{p}(u)+(1_{M(B)}-p)$ for all $u\in\mathcal{U}(\mathcal{M}(A))$. Note that for any $a\in A$ and $u\in\mathcal{U}(\mathcal{M}(A))$ one has that $\phi(ua)=\phi^{\dagger}(u)\phi(a)$.
@@
Let $\phi:A\rightarrow\mathcal{M}(B)$, $\psi:B\rightarrow\mathcal{M}(C)$ be $*{}^{*}$-homomorphisms with $\psi$ extendible, and $q\in\mathcal{M}(C)$ be the corresponding projection to $\psi$. One may compose them by the $*{}^{*}$-homomorphism $\psi_{q}\circ\phi:A\rightarrow\mathcal{M}(C)$ that we simply denote by $\psi\circ\phi$ (see [40, Remark 1.9]). If $\phi$ is also extendible, then the composition $\psi\circ\phi$ remains extendible and $(\psi\circ\phi)^{\dagger}=\psi^{\dagger}\circ\phi^{\dagger}$.
The proposition below yields a more explicit form for the tensor products of bimodules arising from $*{}^{*}$-homomorphisms into the multiplier algebra as in Example 1.3. If the homomorphisms are non-degenerate, then the result is folklore (see [4]).
@@
Suppose that $A,B$, and $C$ are $\mathrm{C}^{*}$-algebras and $\phi:A\to\mathcal{M}(B)$, $\psi:B\to\mathcal{M}(C)$ are $*{}^{*}$-homomorphisms with $\psi$ extendible, but possibly degenerate. Then ${}_{\phi}B\boxtimes{}_{\psi}C\cong{}_{\psi\circ\phi}\overline{\psi(B)C}={}_{\psi\circ\phi}\psi(B)C$.

### 08 Apr 2026 v3

@@ Abstract.
We introduce a categorical approach to classifying actions of C-tensor categories on $\mathrm{C}^{*}$-algebras up to cocycle conjugacy. We show that, in this category, inductive limits exist and there is a natural notion of approximate unitary equivalence. Then, we generalise classical Elliott intertwining results to the tensor category equivariant case, in the same fashion as done by Szabó for the group equivariant case.
RN funded by the European Union. Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the EU nor the ERC can be held responsible for them.
@@ Introduction
Later, Popa classified amenable subfactors $N\subset\mathcal{R}$ by their standard invariant ([38]). Following subsequent reformulations ([31, 32]), the standard invariant of a finite index subfactor $N\subset\mathcal{R}$ can be understood as a pair $(F,Q)$ with $F$ an action of a unitary tensor category $\mathcal{C}$ on NN and $Q$ a special object in $\mathcal{C}$ that allows one to recover the inclusion $N\subset\mathcal{R}$ as a generalised crossed product. [^1] Precisely, $Q$ is a $Q$-system as defined by Longo in [31]. The general setting of $\mathrm{C}^{*}$-tensor categories represents a unifying framework for studying group actions and more general quantum symmetries arising from subfactor theory. Following the recent spectacular classification results for group actions on Kirchberg algebras by group equivariant $K$-theory ([17]), and the development of a $\mathcal{C}$-equivariant $K$-theory ([1]), it is natural to consider to what extent one can classify actions of more general C-tensor categories on simple, amenable C-algebras.
@@
Prior and closely linked to the classification of symmetries of operator algebras was the classification of the operator algebras themselves. First, Connes classified amenable factors in [10] with the exception of one type which was later completed by Haagerup ([23]). In the case of C-algebras, Elliott classified C-inductive limits of finite dimensional C-algebras by their ordered K-theory in [11]. Elliott’s methods laid the foundations for a roadmap to a classification of more general simple, amenable C-algebras. Following Elliott’s strategy and building on decades of work by many mathematicians, the classification programme of simple, amenable $\mathrm{C}^{*}$-algebras successfully culminated in [28, 37, 21, 22, 42, 6].
@@
The general strategy for a classification of operator algebras is to first achieve existence and uniqueness results for morphisms with respect to the proposed classification invariant. Loosely speaking, to construct an isomorphism between two objects AA and $B$ in the category of $\mathrm{C}^{*}$-algebras, by virtue of an existence type statement, one obtains morphisms $\phi:A\to B$ and $\psi:B\to A$ which induce mutually inverse elements at the level of the chosen abstract invariant. Then, a suitable uniqueness result will give that $\psi\circ\phi$ is equivalent to $\mathrm{id}_{A}$ and $\phi\circ\psi$ is equivalent to $\mathrm{id}_{B}$. In the category of separable $\mathrm{C}^{*}$-algebras, the required notion of equivalence is approximate unitary equivalence. The final step is to perform a so-called approximate intertwining argument. Essentially, one can tweak the morphisms $\phi$ and $\psi$ by unitaries until they become mutually inverse isomorphisms (see [40, Corollary 2.3.4] and the introduction of [41] for a more detailed breakdown of this roadmap to classification). Furthermore, in [13], Elliott proposes a similar classification strategy for arbitrary objects in more general categories that have a suitable notion of inner automorphisms.
@@
Instances of the implementation of Elliott’s strategy for the classification of compact (quantum) group actions on C-algebras appear in [19, 2], where the actions are assumed to have the Rokhlin property. In [41], Szabó articulates Elliott’s proposed strategy in the generality of actions of locally compact groups as an alternative to the Evans–Kishimoto intertwining type arguments (see [14]). In his work, Szabó defines an appropriate category of $\Gamma$-C-algebras, that is C-algebras carrying an action of $\Gamma$, and produces a vast collection of intertwining results for this category. Szabó’s construction is a key ingredient facilitating the recent groundbreaking classification of amenable actions on Kirchberg algebras in [17].
@@
In this paper, we will develop the necessary techniques to perform approximate intertwining arguments for C-algebras carrying an action of a tensor category $\mathcal{C}$ ($\mathcal{C}$-$\mathrm{C}^{*}$-algebras). Glimpses of these techniques can be found in [7], where exact intertwining arguments are performed to classify inductive limit actions of fusion categories on AF-algebras. However, for classification, it is often necessary to develop a more general framework which allows to perform approximate intertwining arguments. In the setting of $\mathrm{C}^{*}$-algebras, Elliott used exact intertwining arguments to classify AF-algebras ([11]), but more refined approximate versions were needed to classify simple $A\mathbb{T}$-algebras in [12]. Approximate intertwining arguments in the setting of actions of tensor category have already appeared in the literature; see [43] for an adaptation of the Evans-Kishimoto intertwining argument introduced in [14].
@@
Primarily, the categorical framework developed by Szabó in [41] provides the conceptual skeleton for our construction. However, unlike in the group action case, a tensor category might not act by automorphisms, so our techniques differ. In [41], Szabó defines a cocycle morphism as a pair consisting of an extendible -homomorphism and a unitary cocycle. In the setting of tensor categories, the action on a $\mathrm{C}^{*}$-algebra is given by a family of bimodules acting compatibly with respect to the tensor product. Therefore, the cocycles will be given by certain bimodule maps satisfying a family of commuting diagrams. Hence, it is not apparent how to adapt Szabó’s arguments.
To perform intertwining techniques, we first need to introduce a category whose objects are $\mathcal{C}$-C-algebras. Multiple notions of morphisms between $\mathcal{C}$-C-algebras have appeared in the literature (see [1, Definition 2.11] or [7, Definition 3.2] for example). Any such notion has a common flavour: Given a C-tensor category $\mathcal{C}$ acting on $\mathrm{C}^{*}$-algebras AA and $B$ via tensor functors $(F,J):\mathcal{C}^{\mathrm{rev}}\to\mathrm{Corr}_{0}(A)$ and $(G,I):\mathcal{C}^{\mathrm{rev}}\to\mathrm{Corr}_{0}(B)$ respectively, a morphism between $(A,F,J)$ and $(B,G,I)$ is given by an AA-$B$-correspondence $E$ and a family of bimodule maps $\mathbbm{v}_{X}:F(X)\boxtimes E\rightarrow E\boxtimes G(X)$ satisfying coherence diagrams. [^2] See Section 2 for our conventions on actions of tensor categories. The intricacies between the different definitions lie in how general correspondences we allow and what structure we expect the maps $\mathbbm{v}_{X}$ to have. We choose to work with correspondences which arise from (possibly degenerate) -homomorphisms $\varphi:A\rightarrow B$ and with (possibly non-adjointable) isometries $\mathbbm{v}_{X}$. We will say this data yields a cocycle morphism and we denote the category consisting of $\mathcal{C}$-$\mathrm{C}^{*}$-algebras and cocycle morphisms by $\mathrm{C}^{*}_{\mathcal{C}}$. When the acting category is $\mathrm{Hilb}(\Gamma)$, and the action factors through automorphisms, we give an explicit formula for the family of bimodule maps $\{\mathbbm{v}_{X}\}_{X\in\mathcal{C}}$ in Example $3.$
@@
In $\mathrm{C}^{*}_{\mathcal{C}}$, we are able to construct inductive limits and define a suitable topology on the space of morphisms. An important challenge in constructing a category of cocycle morphisms is the composition. Composing two cocycle morphisms might not give a cocycle morphism. Therefore, we need to introduce a slightly different composition which will allow us to obtain a category. If the morphisms are non-degenerate, this new composition agrees with the canonical one, and with the one introduced by Szabó in the group action case ([41, Proposition 1.15]). Moreover, to avoid the need of restricting to non-degenerate morphisms, the maps $\mathbbm{v}_{X}$ are assumed to be isometries. This assumption allows us, unlike in [41], to work in the possibly non-extendible setting, generalising the notion of a cocycle morphism from [41] when restricting to twisted group actions.
@@
An essential observation for our construction is that we can encode the information of a cocycle morphism into a family of linear maps satisfying certain conditions (see [7, Lemma 3.8] where this is done for unital injective cocycle morphisms). Therefore, a cocycle morphism can be equivalently defined as follows.
@@
Let $\mathcal{C}$ be a $\mathrm{C}^{*}$-tensor category acting on $\mathrm{C}^{*}$-algebras AA and $B$ via $(A,F,J)$ and $(B,G,I)$ respectively. Then a cocycle morphism $(\phi,h):(A,F,J)\rightarrow(B,G,I)$ is given by a -homomorphism $\phi:A\to B$ and a family of linear maps:
@@
    the following diagram commutes:
    $F(Y)\boxtimes F(X)$$F(X\otimes Y)$$G(Y)\boxtimes G(X)$$G(X\otimes Y);$$\scriptstyle{h^{Y}\boxtimes h^{X}}$$\scriptstyle{J_{X,Y}}$$\scriptstyle{h^{X\otimes Y}}$$\scriptstyle{I_{X,Y}}$
@@
In the case of a group action, we give an explicit formula for the family of linear maps $\{h^{X}\}_{X\in\mathcal{C}}$ in Example 3.14. Compared with the usual definition of a cocycle morphism given by pairs $(\phi,\mathbbm{v})$ consisting of a -homomorphism and a family of bimodule maps $\{\mathbbm{v}_{X}\}_{X\in\mathcal{C}}$ (see Definition 3.1), Definition A provides a more direct framework for setting up the intertwining arguments. In particular, to compose two cocycle morphisms, one composes the -homomorphisms, as well as the corresponding linear maps. Furthermore, the natural topology on cocycle morphisms is obtained by considering pointwise differences of the linear maps. When the acting category is semisimple and has countably many isomorphism classes of simple objects (which is assumed for the rest of the introduction), we can phrase convergence in this topology purely in terms of the linear maps.
@@
In general, this topology is coarser than the one used by Szabó in the group action case ([41, Definition 2.5]) (see Example 5.5). However, it is the same when restricted to non-degenerate cocycle morphisms. With this topology in hand, we can then define approximate unitary conjugation for cocycle morphisms. Moreover, unlike Szabó’s definition in [41, Definition 2.8], this is a symmetric relation (see Lemma 5.9), which simplifies the intertwining arguments. This can be formulated as follows.
@@
Along with Theorem D, in Section 6 we use our construction to obtain various $\mathcal{C}$-equivariant versions of classical intertwining arguments. For instance, we establish a general Elliott two-sided intertwining argument (see [40, Proposition 2.3.2] for a non-equivariant version) as well as an intertwining through reparametrisation in Theorem 6.8. These results are essential ingredients in our upcoming classification results for actions of unitary tensor category on Kirchberg algebras in joint work with Kitamura.
In Section 7, the techniques we develop allow us to perfom a $\mathcal{C}$-equivariant one-sided intertwining argument (see Theorem 7.1), generalising the classical one-sided intertwining arguments of [40, Proposition 2.3.5] and [41, Proposition 4.3]. This result is a key ingredient in the work of Evington, C. Jones, and the first named author in obtaining a McDuff-type characterisation of equivariant $\mathcal{D}$-stability of an action of a unitary tensor category, for strongly self-absorbing $\mathcal{D}$ ([15]). The remaining part of Section 7 is concerned with asymptotic versions of the results in Section 6.
@@ 1. Preliminaries
If $X$ only satisfies the properties (i)-(iv) above, then we say $X$ is a pre-Hilbert $B$-module. For $x\in X$, we denote by $|x|^{2}$ the positive element of $B$ given by $\langle x,x\rangle_{B}$. We denote the adjointable operators of $X$ by $\mathcal{L}(X)$. A (right-)Hilbert AA-$B$-bimodule, is a right Hilbert $B$-module with a left AA action by adjointable operators. For a Hilbert AA-$B$-bimodule $X$, we often write the left action of an element $a\in A$ on a vector $x\in X$ by $a\rhd x$ and the right action of an element $b\in B$ on $x$ by $x\lhd b$. The morphisms we consider between Hilbert AA-$B$-bimodules are adjointable linear maps which commute with the left AA-action. In particular, Hilbert AA-$B$-bimodules are called isomorphic if there is a unitary bimodule map between them.
@@
For a C-algebra $B$, the vector space $X=B$ is a Hilbert $B$-module with the right $B$-module structure given by multiplication and the inner product $\langle a,b\rangle_{B}=a^{*}b$ for $a,b\in B$. In this case, any -homomorphism $\phi:A\rightarrow M(B)$ induces a right-Hilbert AA-$B$-bimodule which we denote by $_{\phi}B$.
Let $A,B,C$ be $\mathrm{C}^{*}$-algebras. If $X$ is a Hilbert AA-$B$-bimodule and $E$ is a Hilbert $B$-$C$-bimodule we may form their internal tensor product $X\boxtimes E$ that is a Hilbert AA-$C$-bimodule. We sketch this construction and refer to [29, Section 4] for details. To perform the internal tensor product one starts by considering the algebraic tensor product of vector spaces $X\odot E$. We identify the elements of the form $x\lhd b\odot y$ with $x\odot b\rhd y$ to form the quotient
@@
$\langle x\boxtimes y,z\boxtimes w\rangle_{C}=\langle y,\langle x,z\rangle_{B}\rhd w\rangle_{C},$
for any $x,z\in X$ and $y,w\in E$. It follows that $V$ equipped with this $C$-action and $C$-valued inner product satisfies (i)-(iv) of Definition $1.$ We produce a Hilbert $C$-module $X\boxtimes E$ by completing $V$ under the norm defined by the inner product. Moreover, one can induce a left AA action on $X\boxtimes E$ through
@@
for all $a\in A$, $x\in X$ and $y\in E$. This equips $X\boxtimes E$ with the structure of a Hilbert AA-$C$-bimodule.
If the bimodules are given by -homomorphisms into the multiplier algebra as in Example 1.2, one may get greater insight into the structure of their tensor product. First we recall the definition of non-degenerate and extendible -homomorphisms.
@@
Let AA and $B$ be $\mathrm{C}^{*}$-algebras and $\phi:A\to\mathcal{M}(B)$ a -homomorphism. Then $\phi$ is said to be non-degenerate if $\phi(A)B$ is dense in $B$.
If $\phi:A\to\mathcal{M}(B)$ is non-degenerate, then $\phi$ extends to a unital -homomorphism $\phi:\mathcal{M}(A)\to\mathcal{M}(B)$ ([29, Proposition 2.5]).
@@
Let AA and $B$ be $\mathrm{C}^{*}$-algebras. A -homomorphism $\phi:A\to\mathcal{M}(B)$ is called extendible if for any increasing approximate unit $e_{\lambda}\in A$ (throughout this paper all approximate units are assumed to be increasing), the net $\phi(e_{\lambda})\in\mathcal{M}(B)$ converges strictly to a projection $p\in\mathcal{M}(B)$. [^5] Any non-degenerate -homomorphism $\phi:A\rightarrow\mathcal{M}(B)$ is extendible as for an approximate unit $e_{\lambda}\in A$ the net $\phi(e_{\lambda})$ converges strictly to $1_{\mathcal{M}(B)}$ (see e.g. [3, II.7.3.8.]). As discussed in [41, Definition 1.7], $\phi$ then factorises through $\mathcal{M}(pBp)\cong p\mathcal{M}(B)p\subseteq\mathcal{M}(B)$, with the -homomorphism $\phi_{p}:A\to\mathcal{M}(pBp)$ now non-degenerate. In this case, following [41], we let $\phi^{\dagger}:\mathcal{U}(\mathcal{M}(A))\to\mathcal{U}(\mathcal{M}(B))$ be the unital group homomorphism defined by $\phi^{\dagger}(u)=\phi_{p}(u)+(1_{\mathcal{M}(B)}-p)$ for all $u\in\mathcal{U}(\mathcal{M}(A))$. Note that for any $a\in A$ and $u\in\mathcal{U}(\mathcal{M}(A))$ one has that $\phi(ua)=\phi^{\dagger}(u)\phi(a)$.
@@
Let $\phi:A\rightarrow\mathcal{M}(B)$, $\psi:B\rightarrow\mathcal{M}(C)$ be -homomorphisms with $\psi$ extendible, and $q\in\mathcal{M}(C)$ be the projection corresponding to $\psi$. One may compose them to get the -homomorphism $\psi_{q}\circ\phi:A\rightarrow\mathcal{M}(C)$ that we simply denote by $\psi\circ\phi$ (see [41, Remark 1.9]). If $\phi$ is also extendible, then the composition $\psi\circ\phi$ remains extendible and $(\psi\circ\phi)^{\dagger}=\psi^{\dagger}\circ\phi^{\dagger}$.
The proposition below yields a more explicit form for tensor products of bimodules arising from -homomorphisms into the multiplier algebra as in Example 1.2. If the homomorphisms are non-degenerate, then the result is folklore (see [5]).
@@
Suppose that $A,B$, and $C$ are $\mathrm{C}^{*}$-algebras and $\phi:A\to\mathcal{M}(B)$, $\psi:B\to\mathcal{M}(C)$ are -homomorphisms with $\psi$ extendible, but possibly degenerate. Then $_{\phi}B\boxtimes{}_{\psi}C\cong{}_{\psi\circ\phi}\overline{\psi(B)C}={}_{\psi\circ\phi}\psi(B)C$.